OpenStudy (anonymous):

21. In rectangle KLMN, KM = 4x + 23, and LN = 39. Find the value of x. mc021-2.jpg (1 point) 5 4 16 3

OpenStudy (anonymous):

hmm it looks like you have to have a username and password to open the link.

OpenStudy (anonymous):

Piece of cake :) all you have to do is realize that the diagonals of a rectangle are equal. i.e. KM=LN

OpenStudy (anonymous):

5 4 16 3

OpenStudy (anonymous):

so would it be 5

OpenStudy (anonymous):

if 4x+23=39, what would x be?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

would x=4

OpenStudy (anonymous):

sorry i was away

OpenStudy (anonymous):

and yes it's 4

OpenStudy (anonymous):

22. In rectangle PQRS, PR = 18x – 25, and QS = x + 825. Find the value of x and the length of each diagonal. mc022-1.jpg (1 point) x = 50; PR = 875; QS = 875 x = 51; PR = 893; QS = 893 x = 50; PR = 437.5; QS = 437.5 x = 25; PR = 850; QS = 850

OpenStudy (anonymous):

same principle, different names. PR=QS

OpenStudy (anonymous):

so would it be x = 50; PR = 437.5; QS = 437.5

OpenStudy (anonymous):

idk because i haven't worked it out. I'll leave that to you

OpenStudy (anonymous):

well I got that but some how it feels wrong